OpenStudy (anonymous):

A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies, how many did they sell each week? They sold ? boxes the first week, ? boxes the second week, and ? boxes the third week.

OpenStudy (anonymous):

okay, x can represent the first week first week= x second week= x+5 third week= 2(x+5)=2x+10 the sum is 431

OpenStudy (anonymous):

431=4x+15 solve for x hint: subtract 15 from both sides put the value of x back into the values of each week double check by adding it up

OpenStudy (australopithecus):

girls sold 5 more boxes the second week they doubled sales of the second weekd for the third week to sell 431 boxes of cookies set x = amount of cookies sold the first week set y = amount of cookies sold the second week second week (x + 5) = y third week (x+5)*2 = 431 solve for x then plug that into the second week equation